OpenStudy (anonymous):

How would I graph on a number line, |x+5| ≥ 2

OpenStudy (amistre64):

ack.... number line :) i should learn to read

OpenStudy (anonymous):

on a number line

OpenStudy (amistre64):



OpenStudy (amistre64):

in other words: (x+5)>= 2 or -x-5 >=2 -5 -5 +5 +5 ---------- ----------- x >=-3 -x >= 7 x <= -7

OpenStudy (amistre64):

well, i meant -3 in the picture ....

OpenStudy (amistre64):



OpenStudy (anonymous):

i kni know u amistre64 u remember

OpenStudy (anonymous):

i kni know u amistre64 u remember

OpenStudy (anonymous):

i kni know u amistre64 u remember